the present value of an annuity of $2000 payable at the start of each month at J12 = 12% for 8 years is:
1
Expert's answer
2020-06-15T17:43:53-0400

annuity due:

A = $2000

i = 0.12 / 12 = 0.01 per month

"n = 8\times 12" month


"PV = A\times\frac{1-(1+i)^{-n}}{i}\times(1+i)=2 000\times\frac{1-(1+(0.01))^{-96}}{0.01}\times(1+(0.01))=137822.05"